Aspell .28.3 alpha

A More Intelligent Ispell

Kevin Atkinson

November 20, 1999

Contents

1 Introduction 4

1.1 Comparison to other spell checker engines . . . . . . . . . . . . . . . . . . . . . . . . 4 1.1.1 Comparison to Ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Getting Started 6 2.1 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2 Obtaining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Helping Out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.5 Compiling & Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.5.1 Generic Install Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.5.2 Upgrading from version .28.2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.5.3 Upgrading from version .27.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.5.4 Upgrading from version .25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5.5 Upgrading from version .24 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5.6 Compiling with Egcs and Gcc 2.8 . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.6 Using as a replacement for Ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.6.1 Notes on the Learning from Users Mistakes feature . . . . . . . . . . . . . . . 9 2.6.2 Emacs and Xemacs Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6.3 LYX Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6.4 VIM Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3 The Aspell utility 12

3.1 As a “ispell -a” replacement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

3.1.1 Format of the Data Stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2 As an independent spell checker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.3 As an utility to manage word lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4 Library Interface 18 4.1 The C++ Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 The C Library Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.3 Notes About Thread Safety . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5 International Support 20

6 How Aspell Works 21

A Changelog 22

B To Do 29 B.1 Things that will be done real soon . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 B.2 Things that need to be done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 B.3 Things that I would like to get done . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 B.4 Things that will be done eventually . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 B.5 Good ideas that are worth implementing . . . . . . . . . . . . . . . . . . . . . . . . . 30 B.6 Things that are not likely to get implemented . . . . . . . . . . . . . . . . . . . . . . 31 B.7 Notes and Status of various items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

B.7.1 Affix Compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 B.7.2 Extremely Large Dictionaries . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 B.7.3 General region skipping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 B.7.4 Word skipping by context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 B.7.5 Tablelise to_soundslike . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 B.7.6 Hidden Markov Model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

C Support for Gcc 2.7.2 34 D Credits 35

E Glossary and References 36

F Copyright 38 F.1 LGPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 F.2 Ispell Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 F.3 The Blitz++ Artistic License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

Chapter 1

Introduction

Aspell is an Open Source spell checker designed to eventually replace Ispell. Its main feature is that it does a much better job of coming up with possible suggestions than Ispell does. In fact recent tests shows that it even does better than Microsoft Word 97’s spell checker in some cases. In addition it has both compile time and run time support for other non English languages. Aspell also doubles as a powerful C++ library with C and Perl interfaces in the works.

1.1 Comparison to other spell checker engines

Aspell Ispell Netscape 4.0 Microsoft Word 97
Open Source x  x
Suggestion Intelligence 9  5  7?  8
Personal part of Suggestions x  x  x
Alternate Dictionaries x  x  ?  ?
International Support x  x  ?  ?

The suggestion Intelligence is based on a small test kernel of misspelled/correct word pairs. Go

to http://metalab.unc.edu/kevina/aspell/test for more info and how you can help contribute

to the test kernel. A ? in the suggestion intelligence means that it is a guess based on personal experience.

If you have any other information you would like to add to this chart please contact me at kevinatk@home.com.

1.1.1 Comparison to Ispell

Features that only Aspell has

• Does a much better job with coming up with suggestions than Ispell does.

1.1. COMPARISON TO OTHER SPELL CHECKER ENGINES 5

• Can learn from users misspellings.

• Is an actual library that others programs can link to instead of having to use it through a pipe.

• Is multiprocess intelligent. When a personal dictionary (or replacement list) is saved it will now first update the list against the dictionary on disk in case another process modified it.

• Support for detachable dictionaries so that more than one aspell class can use the same dictionary.

• Support for multiple personal dictionaries as well as support for special auxiliary dictionaries.

Things that only Aspell will have real soon Things that, currently, only Ispell have

• Lower memory footprint

• Support for affix compression

• Support for spell checking Latex and Nroff files.

Getting Started

2.1 Requirements

Aspell requires gcc 2.8.1, egcs, or gcc 2.95 (or better), as the C++ compiler or a reasonable C++ standard compliment compiler with SGI’s STL or STLPort. Unless you are using version 3.11 of SGI’s STL you will need to apply a patch to fix a bug in the rope container. (See secion 2.5.6 for what to do) That latest version of SGI’s STL (http://www.sgi.com/Technology/STL/), version 3.11, fixes the bugs in rope and thus does not need to be fixed.

2.2 Obtaining

The latest version can always be found at Aspell’s home page at http://metalab.unc.edu/kevina/

aspell.

2.3 Support

Support for Aspell can be found on the Aspell mailing list. Instructions for joining the mailing list (and an archive of the mailing list) can be found off the Aspell home page at http://metalab.unc.

edu/kevina/aspell.

2.4 Helping Out

The easiest thing you can do to help out it is too send me your .aspell.ńlangż.rpl file located in your home directory every so often. (Email kevinatk@home.com) This file is used to store the replacement pairs. Unless you specifically use programs that take advantage of this the learning from users mistake feature this file will be empty so there is no point in sending it.

I will then merge your misspelling with the rest of the data in the test kernel which will be used to help improve Aspell’s suggestion intelligence.

Another thing you can do to help out is to try Aspell out on a different language as the Language support system is completely untested.

2.5 Compiling & Installing

2.5.1 Generic Install Instructions

After you have read the sections below and have taken care of any special requirements for you system simply type

./configure && make

or

./configure --disable-static && make

to avoid making the static libraries on a system that supports shared libraries. For additional configure options type ./configure --help.

And Aspell should compile with out any additional user intervention. If you run into problems please send me a note at kevinatk@home.com with the compiler, system you are using and any error messages that were produced.

To install the program simply type

make install

And thats all there is too it.

If something goes wrong first see section 2.5.6 as there is a good change that that will fix your problem.

2.5.2 Upgrading from version .28.2.1

The behavior of “aspell check” changed so that it will now over right the original file as creating new file was creating too many problems when used with programs like pine and vi.

2.5.3 Upgrading from version .27.2

The name of the personal word lists have changes from .aspell.per and .aspell.rpl to .aspell.ńlangż.per and .aspell.ńlangż.rpl respectively. ńlangż is is the language name which will generally be “english”. If you wish to use your old word lists you will need to rename those files.

2.5.4 Upgrading from version .25

The format of the personal replacement dictionary has changed. So, you will either need to rename or remove the file .aspell.rpl located in your home directory. If you have information in this file you would like to preserve please send me an email.

2.5.5 Upgrading from version .24

Because the location of the main word list moved you should probably do a make uninstall (with the old version of a Aspell) before upgrading to remove the old word lists. A make uninstall will not remove any personal word lists.

2.5.6 Compiling with Egcs and Gcc 2.8

In order for aspell to function properly under egcs or gcc 2.8.* you will need to apply the patch file stl_rope-30.diff. (Gcc 2.95 uses version 3.11 of SGI’s STL and thus does not need this patch applied.) To apply the patch cd over to the g++ include directory (normally a subdirectory in the normal include directory) and type in the command

patch < path/misc/stl_rope-30.diff

(where path is the full patch to the location where aspell was unpacked) to patch the necessary file. If it is impossible for you to do this copy the file stl_rope.h to the the directory src/ in the aspell distribution and then type the command:

patch < ../misc/stl_rope-30.diff

Then proceed as normal.

If you do not apply this command aspell will crash when used interactively to spell check files. “aspell -a” and all other aspell modes will work fine however.

2.6 Using as a replacement for Ispell

Aspell can currently function as a drop in replacement for Ispell for programs that use Ispell through a pipe such as Emacs and LYX. It support the basic features of Ispell however it does not currently have a Latex or Nroff mode so there may be situations in which you still wish to use Ispell. Nevertheless, I have been using Aspell for Xemacs and LYX since the middle of September of 1998 with out any problems.

The recommended way to use Aspell as a replacement for ispell is to change the Ispell command from within the program being uses. If that is impossible than the run-with-aspell script can be used. The format of the script is:

run-with-aspell ńcommandż

where ńcommandż is the name of the program with any optional arguments.

The old method of mapping Ispell to Aspell is discouraged because it can create compatibility problems with programs that actually require Ispell such as Ispell’s own scripts.

2.6.1 Notes on the Learning from Users Mistakes feature

Because Ispell does not have this feature most all programs that expect Ispell as the spell checker will not be able to take advantage of this because they do not communicate back the replacement pairs.

If you want to add support for this feature in your program the best way to do it is to scan the version string for the string “Aspell” and if found use the “$$ra” command to communicate back to Aspell the replacement pairs. See section 3.1.1 for more information.

2.6.2 Emacs and Xemacs Notes

The easiest way to use Aspell with Emacs or Xemacs is to add this line:

     (setq-default ispell-program-name "aspell")

to the end of your .emacs file.

For some reason version 3.0 of ispell.el (the lisp program that (x)emacs uses) want to reverse the suggestion list. To fix this add this line:

     (setq-default ispell-extra-args '("--reverse"))

after the previous line in your .emacs file and it should solve the problem.

The latest version of ispell.el, version 3.1 (December 1, 1998), has the list reversing problem fixed.

You can find it at http://www.kdstevens.com/~stevens/ispell-page.html.

2.6.3 LYX Notes

Version 1.0 of LYX provides support for Aspell learning for users mistake feature.

To use aspell with LYX 1.0 either change the spell_command option in the lyxrc file or use the run-with-aspell utility.

2.6.4 VIM Notes

( Note: Due to the change in beahvior of “aspell check” the following will not work without modifiction.)

The following section was written by "R. Marc" <rmarc@copacetic.net>.

Note: The following will not work with Vi as it relies on some Vim specific features.

Impetus: I wanted to use aspell as a replacement for ispell with mutt. I use vim as my default editor in mutt. The macros for ispell listed in the vim faq didn’t quite work the way I wanted them to when replaced with aspell and so the following macro/functions were written.

Macro’s and functions can be place in any file you want (see the vim doc’s for alternative methods if you don’t like mine). I only have a couple macro’s and functions that I’ve defined, so I keep all of it in my .vimrc file.

The order in which you define these makes little difference, so I’ll start with the macro’s. I’ve defined 4 macros for the sake of readability, though you can do them all in one macro if it makes you happy.

map aspell to ˆ T

map \1 :let rm = Aspell_Write()<CR> map \2 :!aspell check %<CR> map \3 :call Aspell_Clean(rm)<CR><CR> map ˆ T \1\2\3

ˆ T is a real control character in this example. In order to add that to your .vimrc using vim you type "<ctrl>-v<ctrl>-t". Macro’s \1 and \3 call user defined vim functions that I’ll define in a moment. The gist of what’s going on is I write the file, spell check it, then write over the existing file with the checked file thus yielding a spell checked document. The way I defined the macro’s has the added benefit of bypassing those silly "Hit RETURN to continue" prompts that Vim is so fond of displaying when various calls are made.

Now about those functions. Here is what they look like:

     function Aspell_Write()
         if strlen(bufname('%'))
             execute ":w!"
         else
             execute ":w! /tmp/Aspellfunc.tmp"
             let rm=1
             return rm
         endif
     endf

     function Aspell_Clean(rm)
         execute ":e! %.checked"
         let tmp=expand("%")
         execute ":file #"
         execute ":w!"
         call delete(expand(tmp))
         if a:rm
             call delete(expand("%"))
         endif
     endf

Aspell_Write checks to see if the file being checked has been named yet or not. If it’s not been named, it writes to a file called /tmp/Aspellfunc.tmp. There is the potential of problems with this if you are spell checking multiple unnamed files at the same time. This isn’t a problem for me, so I didn’t go to great lengths to insure uniqueness. Also, if the file is unnamed, the rm flag is set so we know we need to remove the tmp file when we are done. If it’s already has a name, the file is simply written.

Aspell_Clean takes one argument; the rm flag. It overwrites the existing file with the spell checked version, sets the filename to what it should be and then removes the checked file. If the rm flag is set, the temporary file written for the unnamed file is removed.

The Aspell utility

The Aspell utility is a multipurpose utility that can function as a “ispell -a” replacement, as an independent spell checker, and as a utility for managing dictionaries. Here is a brief summary of Aspell’s command line options.

aspell [options] ńcommandż

ńcommandż is one of:

check ńfileż to check a file pipe ’ispell -a’ compatible mode. dump config dump the current configuration to stdout dump|create|merge maser|personal|repl [word list] dumps, creates, or merges a master, per-

sonal, or replacement word list.

[options] is any or all of the following standard aspell library options:

-d, --master=ńbaseż main word list base name -p, --personal=ńfileż personal word list file name --repl=ńfileż replacements list file name --lang=ńlangż default language to use --dict-dir=ńdirż location of the main word list --data-dir=ńdirż location of language data files --[dont-]ignore-repl [don’t] ignore commands to store replacement pairs

--conf=ńfileż main configuration file

--conf-dir=ńdirż location of main configuration file

in addition to

--reverse reverse the order of the suggestions list.

-b --backup create a backup file by appending “.bak” to the file name. (Only applies when the command is check)

-x --dont-backup don’t create a backup file

In addition Aspell with try to make seance out of Ispell’s command line options so that it can function as a drop in replacement for Ispell when used is “-a” mode.

If Aspell is specified with out any command line options it will display a brief help screen and quit.

Aspell can also make use of a global or user configuration file. Each line of the configuration file has the format:

     ńoptionż [args]

where option is any one of the standard library options above without the leading dashes. For example the following line will set the default language to German:

     lang german

Anything from a “#” to a newline is ignored. The global configuration file is usually named “aspell.conf” and is found in the etc directory while the user configuration file is usually named “.aspell.conf” and is found in the users home directory. Use “aspell dump config” to found out what they are for your installation.

The environmental variable ASPELL_CONF may also be used and it overrides any options set in the configuration file. The format of the string is exactly the same as the configuration file except that semicolons ( ; ) are used instead of newlines.

3.1 As a “ispell -a” replacement

To actually use Aspell as an Ispell replacement simply follow the directions in section 2.6.

When given the pipe or -a command aspell goes into a pipe mode the is compatible with ispell with a few extra additions.

3.1.1 Format of the Data Stream

In this mode, Aspell prints a one-line version identification message, and then begins reading lines of input. For each input line, a single line is written to the standard output for each word checked for spelling on the line. If the word was found in the main dictionary, or your personal dictionary, then the line contains only a ’*’.

If the word is not in the dictionary, but there are suggestions, then the line contains an ’&’, a space, the misspelled word, a space, the number of near misses, the number of characters between the beginning of the line and the beginning of the misspelled word, a colon, another space, and a list of the suggestions separated by commas and spaces.

Finally, if the word does not appear in the dictionary, and there are no suggestions, then the line contains a ’#’, a space, the misspelled word, a space, and the character offset from the beginning of the line. Each sentence of text input is terminated with an additional blank line, indicating that ispell has completed processing the input line.

These output lines can be summarized as follows:

OK: *

Suggestions: & ńoriginalż ńcountż ńoffsetż: ńmissż, ńmissż, ...

None: # ńoriginalż ńoffsetż

When in the -a mode, Aspell will also accept lines of single words prefixed with any of ’*’, ’&’, ’@’, ’+’, ’-’, ’˜ ’, ’#’, ’!’, ’%’, or ’ˆ ’. A line starting with ’*’ tells ispell to insert the word into the user’s dictionary. A line starting with ’&’ tells ispell to insert an all-lowercase version of the word into the user’s dictionary. A line starting with ’@’ causes ispell to accept this word in the future. A line starting with ’+’, ’-’, or ’˜ ’, is ignored for ispell compatibility. A line prefixed with ’#’ will cause the personal dictionary to be saved. A line prefixed with ’!’ will turn on terse mode (see below), and a line prefixed with ’%’ will return ispell to normal (non-terse) mode. Any input following the prefix characters ’+’, ’-’, ’#’, ’!’, ’˜ ’, or ’%’ is ignored, as is any input following. To allow spell-checking of lines beginning with these characters, a line starting with ’ˆ ’ has that character removed before it is passed to the spell-checking code. It is recommended that programmatic interfaces prefix every data line with an uparrow to protect themselves against future changes in Aspell.

To summarize these:

* ńwordż Add a word to the personal dictionary

& ńwordż Insert the all-lowercase version of the word in the personal dictionary

@ ńwordż Accept the word, but leave it out of the dictionary

# Save the current personal dictionary 3.1. AS A “ISPELL -A” REPLACEMENT 15

˜ Ignored + Ignored

- Ignored ! Enter terse mode % Exit terse mode ˆ Spell-check the rest of the line

In terse mode, Aspell will not print lines beginning with ’*’, which indicate correct words. This significantly improves running speed when the driving program is going to ignore correct words anyway. In addition to the above commands which are designed for Ispell compatibility Aspell also supports its own extension. All Aspell extensions follow the following format.

$$ńcommandż [data]

Where data may or may not be required depending on the particular command. Aspell currently supports the following command.

m which Print out the current mode m none Do not use any special mode. m url Enter email, URL, and host name skipping mode. s ńword1ż,ńword2ż Returns the score of the two words based roughly on how aspell would score

them. Sw ńwordż Returns the soundlike equivalent of the word. Sl ńwordż Returns a list of words that have the same soundlike equivalent. Pw ńwordż Returns the phoneme equivalent of the word. pp Returns a list of all words in the current personal wordlist. ps Returns a list of all words in the current session dictionary. l Returns the current language name. ra ńmisż,ńcorż Add the word pair to the replacement dictionary for latter use. Returns nothing. ric Returns the status of the ignore_replacements flag which will either be a 1 or a 0.

3.2. AS AN INDEPENDENT SPELL CHECKER 16 ri0 Sets the ignore_replacement flag to false (the default). Returns nothing. ri1 Sets the ignore_replacement flag to true. Returns nothing. Anything returned is returned on its own line line. All lists returned have the following format

ńnum of itemsż: ńitem1ż, ńitem2ż, ńetcż (Note: part of this section was directly copied out of the Ispell manual) Notes of Storing Replacement Pairs As of version .27 of Aspell storing replacements pairs has a memory. Which means if you first store the replacement pair:

sicolagest -> psycolagest then store the replacement pair

psycolagest -> psychologist The replacement pair

sicolagest -> psychologist will also get stored so that you don’t have to worry about it.

3.2 As an independent spell checker

To use Aspell as an independent spell checker type

aspell check ńfilenameż Where ńfilenameż is the file you want to check. Aspell will create a new file named ńfilenameż.checked with the corrections and the original file is untouched.

3.3 As an utility to manage word lists

To create the main word list from a list of words use the command

aspell --lang=ńlangż create master ./ńbaseż < ńwordlistż

where ńbaseż is the name of the word list and ńword listż is the list of words seperated by white space. The “./” is important because without it aspell will create the word list in the normal word list directory.

This will create two files in the current directory. To use the new word list copy the files to the normal word list directory (use “aspell dump config” to find out what it is) and use the option --master=ńbaseż.

A personal and replacement word list can be created in a similar fashion.

Please note that Aspell currently does not support any sort of affix compression like Ispell does. Thus Ispell’s word list will not work as is. In order to use Ispell’s word lists simply pipe the word list through “ispell -e” to expand the munched word lists.

The replacement word has each replacement pair on its own line in the following format

     ńmisspelled wordż: ńcorrectionż

The dump command will simply dump the contents of a word list to stdout in a formate than can be read back in with aspell create.

If no word list is specified the command will act on the default one. For example the command

aspell dump master

will simply dump the contents of the current master word list to stdout.

Library Interface

4.1 The C++ Library

The new C++ library makes heavy use of modern C++ features and the STL. Documenation should be completed soon. For right now please see the header files installed with the distribution (usually under /usr/local/include/aspell/). For examples of how to use the C++ please see the source for the aspell utility (located under src/aspell.cc)

4.2 The C Library Interface

I am also providing a C interface so that C programmers can use my library without a lot of hassle. Unfortunately because my actually library is C++ code there are several cravats to linking C to C++ code according to the C++FAQ Lite:

• Your must use your C++ compiler when compiling main() (e.g., for static initialization)

• Your C++ compiler should direct the linking process (e.g., so it can get its special libraries)

• Your C and C++ compilers probably need to come from same vendor and have compatible versions (e.g., so they have the same calling conventions)

However depending on the compiler and/or linker you are using you may be able to get away with not doing one or more of these things. Your mileage may vary.

Unfortunately due to a major rewrite of the C++ library the C library is currently unavailable. I hope to have it back in a release or two.

4.3 Notes About Thread Safety

Readonly Aspell methods and functions should be thread safe as long as exceptions, new, delete, delete[], and STL allocators are thread safe. To the best of my knowledge gcc and egcs meet these requirements. It is up to the programmer to make sure multiple threads do not do thing such as change the dictionaries and add or delete items from the personal or session dictionaries.

International Support

Aspell international support is currently under going a major rewrite. So, for right now the best thing to do if you want to see your language supported is to send me the following information:

• The iso-* character set used for your language

• Any nonword characters that can appear in a word such as a dash

• Any other special considerations for you language such as allowing run together words.

• A small word list of around 1,000 words so that I can test things out.

I will then get back to you with necessary files and instructions on how to get things set up. My email is kevinatk@home.com.

More information on my future planes for international support in aspell can be found at http:

//metalab.unc.edu/kevina/aspell/international/.

How Aspell Works

The magic behind my spell checker comes from merging Lawrence Philips excellent metaphone algorithm and Ispell’s near miss strategy which is inserting a space or hyphen, interchanging two adjacent letters, changing one letter, deleting a letter, or adding a letter.

The process goes something like this.

1. Convert the misspelled word to its soundslike equivalent (its metaphone for English words).

2. Find words that have the same soundslike pattern.

3. Find words that have similar soundslike patterns. A similar soundlike pattern is a pattern that is obtained by interchanging two adjacent letters, changing one letter, deleting a letter, or adding a letter.

4. FInd misspelled words that have a correctly spelled replacement by the same criteria of step number 2 and 3. That is the misspelled word in the word pair (such as teh->the) would appear in the suggestions list as if it was a correct spelling.

5. Score the result list and return the words with the lowest score. The score is roughly the weighed average of the edit distance of the word to the misspelled word, the soundslike equivalent of the two words, and the phoneme of the two words. The edit distance is the weighed total of the number of deletions, insertions, exchanges, or adjacent swaps needed to make one string equivalent to the other.

6. Replace the misspelled words that have correctly spelled replacements with their replacements and remove any duplicates that might arise because of this.

Please note that the soundslike equivalent is a rough approximation of how the words sounds. It is not the phoneme of the word by any means. For more details about exactly how each step is performed please see the file suggest.cc. For more information on the metaphone algorithm please see the file metaphone.cc which included a detailed description of it.

Appendix A

Changelog

Changes from .28.2.1 to .28.3 (Nov 20, 1999)

• Fixed a bug that caused aspell to crash when spell checking words over 60 characters long.

• Reworked “aspell check” so that 1. You no longer have to hit enter when making a choice.

2. It will now overwrite the orignal file instead of creating a new file. An optional backup

can be made by using the -b option.

• Fixed a few bugs in data.cc.

Changes from .28.2 to .28.2.1 (Aug 25, 1999)

• Fixed the version number for the shared library.

• Fixed a problem with undefined refrences when linking to the shared library.

Changes from .28.1 to .28.2 (Aug 25, 1999)

• Fixed a bunch of bugs in the language and configuration classes.

• Minor changed in the code so that it could compile with the new gcc 2.95(.1).

• Changed the output of “dump config” so that default values are given the value "<default>". This means that the output can be used to created a configuration file.

• Added notes on using aspell with VIM.

Changes from .28 to .28.1 (July 27, 1999)

• Removed some debug output

• Changed notes on compiling with gcc 2.8.* as I managed to get it to compile on my school account

• Avoided included stdexcept in const_string.hh so that I could get to compile on my schools account with gcc 2.8.1.

Changes from .27.2 to .28 (July 25, 1999)

• Provided an iterator for the replacement classes.

• Added support for dumping and creating and merging the personal and replacement word lists.

• Changed the aspell utility command line a bit, it now used popt.

• Totally reworked aspell configuration system. Now aspell can get configuration from any of 5 sources: the command line, the environmental variable ASPELL_CONF, the personal configuration file, the global configuration file, and finally the compiled in defaults.

• Totally reworked the language class in preparation for my new language code. See http:

  //metalab.unc.edu/kevina/aspell/international/ for more information of what I have

in store.

• Added some options to the configure script: --enable-dict-dir=DIR, --enable-doc-dir=DIR, --enable-debug, and --enable-opt

• Removed some old header files.

• Reoriginized the directory structure a bit

• Made the text version of the manual pages slightly easier to read

• Used the \url command for urls for better formating of the printed version.

Changes from .27.1 to .27.2 (Mar 1, 1999)

• Fixed a major bug that caused aspell to dump core when used without any arguments

• Fixed another major bug that caused aspell to do nothing when used in interactive mode.

• Added an option to exit in Aspell’s interactive mode.

• Removed some old documentation files from the distribution.

• Minor changes on to the section on using Aspell with egcs.

• Minor changes to remove -Wall warnings.

Changes from .27 to .27.1 (Feb 24, 1999)

• Fixed a minor compile problem.

• Updated the section on using Aspell with egcs to it it more clear why the patch is necessary.

Changes from .26.2 to .27 (Feb 22, 1999)

• Totally reworked the C++ library which means you may need to change some things in your code.

• Added support for detachable and multiple personal dictionaries in the C++ class library.

• The C++ class library now throws exceptions.

• Reworked aspell ability to learn from users misspellings a bit so that it now has a memory. See section 3.1.1 for more information.

• Upgraded autoconf to version 2.13 and automake to version 1.4 for better portability.

• Fixed the configuration so the “make dist” will work. For now on aspell will be distributed with “make dist”.

• Added support to skip over URL’s, email addresses and host names.

• Added support for dumping the master and personal word list. You can now also merge a personal word list. Type aspell -help for help on using this feature.

• Reorganized the source code.

• Started using proper version numbers for the shared library.

• Fixed a bug that caused a spell to crash when adding certain replacement pairs.

• Fixed the problem with duplicate lines when exiting pipe mode for good.

Changed from .26.1 to .26.2 (Jan 3, 1998)

• Fixed another compile problem. Hopefully this time it will really compile OK on other peoples machines.

Changed from .26 to .26.1 (Jan 3, 1998)

• Fixed a small compile problem in as_data.cc.

Changed from .25.1 to .26 (Jan 3, 1999)

• Fixed a bug that causes duplicates items to be displayed in the suggestion list for good. (If it still does it please send be email.)

• Added the ability for aspell to learn form the users misspellings.

• Library Interface changes. Still more to come....

• Is now multiprocess safe. When a personal dictionary (or replacement list) is saved it will now first update he list against the dictionary on disk in case another process modified it.

• Fixed the bug that caused duplicate output when used non interactively in pipe mode.

• Dropped support for gcc 2.7.2 as the C++ compiler.

• Updated the How Aspell Works (6) Chapter.

• Added support for the ASPELL_DATA_DIR environmental variable.

Changes from .25 to .25.1 (Dec 10, 1998)

• Fixed the version number so that Aspell reports the correct version number.

• Changed the note on gcc 2.7.2 compilers to make it clear that only the C++ compiler can not be gcc 2.7.2, it is ok if the C compiler is gcc 2.7.2.

• Updated the TODO list and reorganized it a bit.

• Fixed the directory so that all the documentation will get installed in ${prefix}/doc/aspell instead of half of it in ${prefix}/doc/aspell and half of it in ${prefix}/doc/kspell.

Changes from .24 to .25 (Nov 23, 1998)

• Total rework of how the main word list is stored. Start up time decreased to about 1/3 of what it was in .24 and memory usage decreased to about 2/3. (When used with the provided word list on a linux system). Also the format and default locations of the main word list data files changed in the process and the data is now machine dependent. The personal word list format, however, stayed the same.

• Changed the scoring method to produce slightly better results with words like the vs. teh. And other simpler misspellings where two letters are swapped.

• Fixed the very unpredictable behavior of the ’*’, ’&’, ’@’ commands in the pipe mode.

• Added documentations for Aspell pipe mode (also known as ispell -a compatibility mode)

• Added a bunch of Aspell specific extensions to the pipe mode and documented them.

• Documented the to_soundslike and soundslike methods for the aspell class.

• Changed the scoring method to produce better results for words like “fone” vs “phone” and other words that have a spelling that doesn’t directly relate to how the word sounds by using the phoneme equivalent of the word in the scoring of it.

• Added the to_phoneme and have_phoneme methods to the SC_Language class.

• Added the to_phoneme method to the aspell class.

• Added the framework for being able to learn from the users misspelling. Right now it just keep a log of replacements.

• Redid stl_rope-30.diff. For some reason the version of patch on my system refused it.

• Rewrite of the “Using as a replacement for Ispell” section and added the run-with-aspell utility as a replacement of the old method of mapping Ispell to Aspell.

• Fixed a bug that caused duplicate words to appear in the suggestion list.

Changes from .23 to .24 (Nov 8, 1998)

• Fixed my code so that it can once again compile with g++ 2.7.2.

• Rewrote the How It Works chapter.

• Rewrote the Requirement section and added noted on compiling with g++ 2.7.2.

• Added a To Do chapter.

• Added a Glossary and References chapter.

• Other minor documentation improvements.

• Internal code documentation improvements.

Changes from .22.1 to .23 (Oct 31, 1998)

• Minor documentation fixes.

• Changed the scoring strategy for words with 3 or less letters. This cut the number of words returned for these roughly in half.

• Expanded the word list to also include american.0 and american.1 from the Ispell distribution. It now includes english.0, english.1, american.0 and american.1 from the directory languages/english provided with Ispell 3.1.20.

• Added a link to the location of the latest Ispell.el in the documentation.

• Started a C interface and added some rough documentation for it.

Changes from .22 to .22.1 (Oct 27, 1998)

• Minor bug fixes. I was deleting arrays with delete rather than delete[]. I was suppressed that this had not created a problem.

• Added a simple test program to test for a memory leak present on some systems. (Only debian slink at the moment.) See the file memleak-test.cc for more info.

Changes from .21 to .22 (Oct 26, 1998)

• Major redesign or the scoring method. It now uses absolute distances rather than relative scores for more consistent results. See suggest.cc for more info.

• Suggest code rewritten is several places however the core process stayed the same.

• The suggest_ultra method temporally does nothing. It should be working again by the next release.

Changes from .20 to .21 (Oct 13, 1998)

• Added documentation for aspell::Error

• Changed the library name from libspell to libaspell. It should never have been libspell in the first place. Sorry for the incompatibility.

• Added as_error.hh to the list of files copied to the include directory so that you can actually use the library outside of the source dir.

• Fixed bug that caused a segmentation fault with words where the only suggestions was inserting a space or hyphen such as in “ledgerline”.

• Added the score method to aspell.

• Changed the scoring method to deal with word when the user uses "f" in place of "ph" a lot better.

Changes from .11 to .20 (Oct 10, 1998)

• Name change. Everything that was Kspell is now Aspell. Sorry, the name Kspell was already used by KDE and I didn’t want to cause any confusion.

• Fixed a bug that causes a segmentation fault when the HOME environmental variable doesn’t exist.

Changes from .10 to .11 (Sep 12, 1998)

• Overhaul of the SC_Language class

• Added documentation for international support

• Added documentation for the C++ library

• Other minor bug fixes.

Appendix B

To Do

Words in bold indicate how you should refer to the item when discussing it with me or others.

B.1 Things that will be done real soon

These items should get done within a release or two.

• Handle capitulation better when storing replacement pairs.

• Finish up support for detachable and multiple dictionaries, right now aspell will not work right with multiple replacement dictionaries.

• Totally rewrite the aspell international support. See http://metalab.unc.edu/kevina/

  aspell/international/ for more information.

• Rework the aspell check function to provide support for using any number of filters which will be needed for international support.

• Namespace clean up. Chose a naming sceme for all of aspell functions and classes and stick to it.

• Write documentation for the library interface.

• Bring back the C interface library.

B.2 Things that need to be done

Things items will eventually be implemented as I know they are important however I am not sure when they will get done.

• Add support for affix compression.

• Figure out a way for Aspell to work better with extremely large dictionaries.

B.3 Things that I would like to get done

These items will eventually be implemented. I hope to have them all done before I move aspell to beta testing. They are in the approximate order of when they will get done.

B.4 Things that will be done eventually

I plan on doing these things eventually. It is just a matter of getting around to it.

• Bring the suggest_ultra method back.

B.5 Good ideas that are worth implementing

These items all sound like good ideas however I am not sure when I will get to implementing then if ever. If you are looking for a way to contribute picking up on one of these ideas would be a great way to start. They are presented in no particular order.

• Come up with a full screen terminal interface similar to Ispell’s.

• Come up with a plug-in for gEdit the gnome text editor.

• Create a debian package for both the Aspell utility and the library.

• Create a RPM package for both the Aspell utility and the library.

• Change languages (and thus dictionaries) based on the information in the actual document.

• Come up with an HTML mode for spell checking.

• Come up with a TEX mode and a nroff mode for spell checking.

• Come up with a C mode, a C++ mode, a Perl mode, etc for spell checking.

• Come up with a mode that will skip words based on the symbols that (almost) always surround the word. (Word skipping by context)

• Create two server modes for Aspell. One that uses the DICT protocol and one that uses ispell -a method of communication of some arbitrary port.

• Improve the C interface and documentation.

• Come up with a Perl interface.

• Come up with thread safe personal dictionaries.

• Tablelise the to_soundslike method.

• Use the Hidden Markov Model to base the suggestions on not only the word itself but on the context around the word.

B.6 Things that are not likely to get implemented

Theses ideas are not likely to get implemented any time soon.

• (None Yet)

B.7 Notes and Status of various items

B.7.1 Affix Compression

Due to the current way my spell checker works implementing affix compression would be next to impossible. Nevertheless, I do realize that for some languages affix compression is very important.

So to solve this dilemma I plan on having two different modes of my spell checker: One with affix compression that does not use soundslike pairs (much like ispell) and one without affix compression that does use soundslike.

I plan to extract the affix manipulation code from Ispell with the help of an Ispell author. The tricky part would be getting this to getting this all to work properly at tun time bases on the dictionary used.

B.7.2 Extremely Large Dictionaries

This problem extends back to the fact of the way words are index is Aspell. This problem will get resolved when I implanted the affix compression mode as only one index would be used.

B.7.3 General region skipping

I want to implement this give other people an idea of how it should be done and because I am really sick of having to spell check through url and email address.

B.7.4 Word skipping by context

This was posted on the Aspell mailing list on January 1, 1999:

I had an idea on a great general way to determine if a word should be skipped. Determine the words to skip based on the symbols that (almost) always surround the word.

For example when asked to check the following C++ code:

     cout << "My age is: " << num << endl;
     cout << "Next year I will be " << num + 1 << endl;

cout, num, and endl will all be skipped. "cout" will be skipped because it is always preceded by a <<. "num" will be skipped because it is always preceded by a <<. And "endl" will be skipped because it is always between a << and a ;.

Given the following html code.

     <table width=50% cellspacing=0 cellpadding=1>
     <tr><td>One<td>Two<td>Three
     <tr><td>1<td>2<td>3
     </table>
     <table cellspacing=0 cellpadding=1>
     </table>

table, width cellspacing, cellpadding, tr, td will all be skipped because they are always inclosed in "<>". Now of course table and width would be marked as correct anyway however there is no harm in skipping them.

So I was wondering if anyone on this list has any experience in writing this sort of context recognition code or could give me some pointers in the right direction.

This sort of word skipping will be very powerful if done right. I imagine that it could replace specific spell checker modes for Tex, Nroff, SGML etc because it will automatically be able to figure out where it should skip words. It could also probably do a very good job on programming languages code.

If you are interested in helping be out with this or just have general comments about the idea please let me know.

B.7.5 Tablelise to_soundslike

Geoff Kuenning (one of the Ispell authors) gave me this idea. He suggest that I table-drive the conversion instead of hard wiring the code to encourage other people to come up with a to_soundslike method for their language. “My experience with ispell leads me to believe that if we could give fairly explicit instructions about how to design a metaphone algorithm for a language, there are people out there who are eager to do the work.”

This sounds like a good idea to and it is something I hope to implement eventually as it would surely beat my current crude method of just deleting all the values for languages without a customized to_soundslike method for the language.

B.7.6 Hidden Markov Model

Knud Haugaard Sørensen suggested this one. From his email on the Aspell mailing list:

consider this examples.

a fone number. -> a phone number.

a fone dress. -> a fine dress.

the example illustrates that the right corection might depend on the context of the word. So I suggest that you take a look on HMM to solve this problem.

This might also provide a good base to include gramma correction in aspell.

     see this link http://www.cse.ogi.edu/CSLU/HLTsurvey/ch1node7.html

I think it is a great idea. However unfortunately it will probably be very complicated to implement. Perhaps in the far future...

Appendix C

Support for Gcc 2.7.2

(and other non-standard compliance compilers)

My original plan was to program in such a way that it would Aspell would compile under Gcc 2.7.2 however after releasing a rather nasty bug in 2.7.2 with nested typed I have desired to drop all support for G++ 2.7.2. As of Aspell .27 all hope for being able to compile under Gcc 2.7.2 is lost as I am now using many modern C++ features which are simply not present in C++ most notably template specialization and template members. Egcs 1.1.1 is a very good standards compliant compiler and that is now the officially supported compiler. However as Egcs 1.1.1 is relatively new and, except for namespaces, provides little new functionality I will continue to support Egcs 1.0.3. Gcc 2.8.1 should in theory work however it is so buggy I have yet to get it to compile with it nor has anyone else that I know of.

Yes, my code could be rewritten so that is could compile under Gcc 2.7.2 but Why? Using modern C++ has probably accelerated the development of this library my at least 50%. And for that matter why stop at Gcc 2.7.2 why not go all out and totally rewrite my code pure C. I hope you see my point.

However that does not mean I want to sacrifice portability unnecessarily. If you see any of part of my code that in not Standard C++ please let me know. My hope is that my code could compile on all Standard compliance C++ compilers with the addition of a few extra header files from SGI’s STL.

As I side note I think that Mozilla’s C++ portability guide (http://www.mozilla.org/docs/ tplist/catBuild/portable-cpp.html) could be summed up in one sentence: Program in the dark ages of C++.

Appendix D

Credits

• To the many authors of Ispell (including R. E. Gorin, Pace Willisson, and Geoff Kuenning) for providing me with a good word list as well as giving me a few good ideas.

• Lawrence Philips for coming up with the original metaphone algorithm and Michael Kuhn for writing C code for the algorithm.

• To the authors of SGI STL version 3.0 and up for proving a great set of generic container classes with cut the development time of this program in half at least.

• To the LYX development team for giving me a great tool for the development of this manual.

• JAmes Atwill for providing me an account on his machine so that I can once again get Aspell to compile with g++ 2.2.2

Appendix E

Glossary and References

affix in grammar, a word element that, when added to a word, modifies its meaning or function;

prefix, infix, or suffix. Debian A 100% Open Source linux distribution http://www.debian.org DICT protocol A TCP transaction based query/response protocol that allows a client to access

dictionary definitions from a set of natural language dictionary databases. http://www.dict.

     org/

Gnome A project to build a complete, user-friendly desktop based entirely on free software. http:

     //www.gnome.org/

GTK+ A library for creating graphical user interfaces for the X Window System. It is designed

to be small, efficient, and flexible. http://www.gtk.org/ GUI Graphics User Interface

Ispell.el Emacs interface for ispell. http://www.kdstevens.com/~stevens/ispell-page.html

Ispell An international spell checker which is just about the only decent Open Source spell checker

     out there. (except Aspell of course). http://fmg-www.cs.ucla.edu/geoff/ispell.html

Ispell -a An Ispell mode that is designed to be used by other applications though a pipe. KDE A powerful graphical desktop environment for Unix workstations. http://www.kde.org Linux A Open Source version of Unix which runs on many platforms. http://www.linux.org/ LYX An What You See is What You Mean document editor for the X environment, http://www.

     lyx.org

Open Source Software where the source code is available for anyone to extend or modify. http:

     //www.opensource.org/

Red Hat A commercial Linux distribution. http://www.redhat.com RPM Red Hat’s packing format also used by other Linux distributions. STL Standard Temple Library. A C++ library of container classes, algorithms, and iterators. SGI STL Silicon Graphics implantation of the STL. http://www.sgi.com/Technology/STL/ STLPort A port of SGI STL designed to run on compilers that don’t support all the latest features

     of C++. http://corp.metabyte.com/~fbp/stl/

Appendix F

Copyright

The library and the utility program is Copyrighted 1998 by Kevin Atkinson under the LGPL for now. However the LGPL is generally not suitable for modern C++ as inline functions and templates tend to make it difficult to draw a line between what is a program and what is a library. I am thinking about using a license smiler to the The Blitz++ Artistic License.

The provided word list is extracted from the word lists provided with Ispell thus it is distributed under the same terms of Ispell.

The three licenses follow:

F.1 LGPL

GNU LIBRARY GENERAL PUBLIC LICENSE Version 2, June 1991

Copyright (C) 1991 Free Software Foundation, Inc.

59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

[This is the first released version of the library GPL. It is numbered 2 because it goes with version 2 of the ordinary GPL.]

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public Licenses are intended to guarantee your freedom to share and change free software–to make sure the software is free for all its users.

This license, the Library General Public License, applies to some specially designated Free Software Foundation software, and to any other libraries whose authors decide to use it. You can use it for your libraries, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the library, or if you modify it.

For example, if you distribute copies of the library, whether gratis or for a fee, you must give the recipients all the rights that we gave you. You must make sure that they, too, receive or can get the source code. If you link a program with the library, you must provide complete object files to the recipients so that they can relink them with the library, after making changes to the library and recompiling it. And you must show them these terms so they know their rights.

Our method of protecting your rights has two steps: (1) copyright the library, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the library.

Also, for each distributor’s protection, we want to make certain that everyone understands that there is no warranty for this free library. If the library is modified by someone else and passed on, we want its recipients to know that what they have is not the original version, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that companies distributing free software will individually obtain patent licenses, thus in effect transforming the program into proprietary software. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

Most GNU software, including some libraries, is covered by the ordinary GNU General Public License, which was designed for utility programs. This license, the GNU Library General Public License, applies to certain designated libraries. This license is quite different from the ordinary one; be sure to read it in full, and don’t assume that anything in it is the same as in the ordinary license.

The reason we have a separate public license for some libraries is that they blur the distinction we usually make between modifying or adding to a program and simply using it. Linking a program with a library, without changing the library, is in some sense simply using the library, and is analogous to running a utility program or application program. However, in a textual and legal sense, the linked executable is a combined work, a derivative of the original library, and the ordinary General Public License treats it as such.

Because of this blurred distinction, using the ordinary General Public License for libraries did not effectively promote software sharing, because most developers did not use the libraries. We concluded that weaker conditions might promote sharing better.

However, unrestricted linking of non-free programs would deprive the users of those programs of all benefit from the free status of the libraries themselves. This Library General Public License is intended to permit developers of non-free programs to use free libraries, while preserving your freedom as a user of such programs to change the free libraries that are incorporated in them. (We have not seen how to achieve this as regards changes in header files, but we have achieved it as regards changes in the actual functions of the Library.) The hope is that this will lead to faster development of free libraries.

The precise terms and conditions for copying, distribution and modification follow. Pay close attention to the difference between a "work based on the library" and a "work that uses the library". The former contains code derived from the library, while the latter only works together with the library.

Note that it is possible for a library to be covered by the ordinary General Public License rather than by this special one.

GNU LIBRARY GENERAL PUBLIC LICENSE TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License Agreement applies to any software library which contains a notice placed by the copyright holder or other authorized party saying it may be distributed under the terms of this Library General Public License (also called "this License"). Each licensee is addressed as "you".

A "library" means a collection of software functions and/or data prepared so as to be conveniently linked with application programs (which use some of those functions and data) to form executables.

The "Library", below, refers to any such software library or work which has been distributed under these terms. A "work based on the Library" means either the Library or any derivative work under copyright law: that is to say, a work containing the Library or a portion of it, either verbatim or with modifications and/or translated straightforwardly into another language. (Hereinafter, translation is included without limitation in the term "modification".)

"Source code" for a work means the preferred form of the work for making modifications to it. For a library, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the library.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running a program using the Library is not restricted, and output from such a program is covered only if its contents constitute a work based on the Library (independent of the use of the Library in a tool for writing it). Whether that is true depends on what the Library does and what the program that uses the Library does. 1. You may copy and distribute verbatim copies of the Library’s complete source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and distribute a copy of this License along with the Library.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Library or any portion of it, thus forming a work based on the Library, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions:

a) The modified work must itself be a software library.

b) You must cause the files modified to carry prominent notices stating that you changed the files and the date of any change.

c) You must cause the whole of the work to be licensed at no charge to all third parties under the terms of this License.

d) If a facility in the modified Library refers to a function or a table of data to be supplied by an application program that uses the facility, other than as an argument passed when the facility is invoked, then you must make a good faith effort to ensure that, in the event an application does not supply such function or table, the facility still operates, and performs whatever part of its purpose remains meaningful.

(For example, a function in a library to compute square roots has a purpose that is entirely welldefined independent of the application. Therefore, Subsection 2d requires that any applicationsupplied function or table used by this function must be optional: if the application does not supply it, the square root function must still compute square roots.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Library, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Library, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Library.

In addition, mere aggregation of another work not based on the Library with the Library (or with a work based on the Library) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may opt to apply the terms of the ordinary GNU General Public License instead of this License to a given copy of the Library. To do this, you must alter all the notices that refer to this License, so that they refer to the ordinary GNU General Public License, version 2, instead of to this License. (If a newer version than version 2 of the ordinary GNU General Public License has appeared, then you can specify that version instead if you wish.) Do not make any other change in these notices.

Once this change is made in a given copy, it is irreversible for that copy, so the ordinary GNU General Public License applies to all subsequent copies and derivative works made from that copy.

This option is useful when you wish to copy part of the code of the Library into a program that is not a library.

4. You may copy and distribute the Library (or a portion or derivative of it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange.

If distribution of object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place satisfies the requirement to distribute the source code, even though third parties are not compelled to copy the source along with the object code.

5. A program that contains no derivative of any portion of the Library, but is designed to work with the Library by being compiled or linked with it, is called a "work that uses the Library". Such a work, in isolation, is not a derivative work of the Library, and therefore falls outside the scope of this License.

However, linking a "work that uses the Library" with the Library creates an executable that is a derivative of the Library (because it contains portions of the Library), rather than a "work that uses the library". The executable is therefore covered by this License. Section 6 states terms for distribution of such executables.

When a "work that uses the Library" uses material from a header file that is part of the Library, the object code for the work may be a derivative work of the Library even though the source code is not. Whether this is true is especially significant if the work can be linked without the Library, or if the work is itself a library. The threshold for this to be true is not precisely defined by law.

If such an object file uses only numerical parameters, data structure layouts and accessors, and small macros and small inline functions (ten lines or less in length), then the use of the object file is unrestricted, regardless of whether it is legally a derivative work. (Executables containing this object code plus portions of the Library will still fall under Section 6.)

Otherwise, if the work is a derivative of the Library, you may distribute the object code for the work under the terms of Section 6. Any executables containing that work also fall under Section 6, whether or not they are linked directly with the Library itself.

6. As an exception to the Sections above, you may also compile or link a "work that uses the Library" with the Library to produce a work containing portions of the Library, and distribute that work under terms of your choice, provided that the terms permit modification of the work for the customer’s own use and reverse engineering for debugging such modifications.

You must give prominent notice with each copy of the work that the Library is used in it and that the Library and its use are covered by this License. You must supply a copy of this License. If the work during execution displays copyright notices, you must include the copyright notice for the Library among them, as well as a reference directing the user to the copy of this License. Also, you must do one of these things:

a) Accompany the work with the complete corresponding machine-readable source code for the Library including whatever changes were used in the work (which must be distributed under Sections 1 and 2 above); and, if the work is an executable linked with the Library, with the complete machine-readable "work that uses the Library", as object code and/or source code, so that the user can modify the Library and then relink to produce a modified executable containing the modified Library. (It is understood that the user who changes the contents of definitions files in the Library will not necessarily be able to recompile the application to use the modified definitions.)

b) Accompany the work with a written offer, valid for at least three years, to give the same user the materials specified in Subsection 6a, above, for a charge no more than the cost of performing this distribution.

c) If distribution of the work is made by offering access to copy from a designated place, offer equivalent access to copy the above specified materials from the same place.

d) Verify that the user has already received a copy of these materials or that you have already sent this user a copy.

For an executable, the required form of the "work that uses the Library" must include any data and utility programs needed for reproducing the executable from it. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

It may happen that this requirement contradicts the license restrictions of other proprietary libraries that do not normally accompany the operating system. Such a contradiction means you cannot use both them and the Library together in an executable that you distribute.

7. You may place library facilities that are a work based on the Library side-by-side in a single library together with other library facilities not covered by this License, and distribute such a combined library, provided that the separate distribution of the work based on the Library and of the other library facilities is otherwise permitted, and provided that you do these two things:

a) Accompany the combined library with a copy of the same work based on the Library, uncombined with any other library facilities. This must be distributed under the terms of the Sections above.

b) Give prominent notice with the combined library of the fact that part of it is a work based on the Library, and explaining where to find the accompanying uncombined form of the same work.

8. You may not copy, modify, sublicense, link with, or distribute the Library except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, link with, or distribute the Library is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

9. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Library or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Library (or any work based on the Library), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Library or works based on it.

10. Each time you redistribute the Library (or any work based on the Library), the recipient automatically receives a license from the original licensor to copy, distribute, link with or modify the Library subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

11. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Library at all. For example, if a patent license would not permit royalty-free redistribution of the Library by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Library.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply, and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

12. If the distribution and/or use of the Library is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Library under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

13. The Free Software Foundation may publish revised and/or new versions of the Library General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Library specifies a version number of this License which applies to it and "any later version", you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Library does not specify a license version number, you may choose any version ever published by the Free Software Foundation.

14. If you wish to incorporate parts of the Library into other free programs whose distribution conditions are incompatible with these, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

15. BECAUSE THE LIBRARY IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE LIBRARY, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE LIBRARY "AS IS" WITHOUT WARRANTY OF ANY KIND, EI-THER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE LIBRARY IS WITH YOU. SHOULD THE LIBRARY PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE LIBRARY AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUEN-TIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE LIBRARY (IN-CLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INAC-CURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE LIBRARY TO OPERATE WITH ANY OTHER SOFTWARE), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Libraries

If you develop a new library, and you want it to be of the greatest possible use to the public, we recommend making it free software that everyone can redistribute and change. You can do so by permitting redistribution under these terms (or, alternatively, under the terms of the ordinary General Public License).

To apply these terms, attach the following notices to the library. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the "copyright" line and a pointer to where the full notice is found.

<one line to give the library’s name and a brief idea of what it does.>

Copyright (C) <year> <name of author>

This library is free software; you can redistribute it and/or modify it under the terms of the GNU Library General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

This library is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FIT-NESS FOR A PARTICULAR PURPOSE. See the GNU Library General Public License for more details.

You should have received a copy of the GNU Library General Public License along with this library; if not, write to the Free Foundation, Inc., 59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Also add information on how to contact you by electronic and paper mail.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a "copyright disclaimer" for the library, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in the library ‘Frob’ (a library for tweaking knobs) written by James Random Hacker.

<signature of Ty Coon>, 1 April 1990 Ty Coon, President of Vice

That’s all there is to it!

F.2 Ispell Copyright

Copyright (c), 1983, by Pace Willisson

Copyright 1992, 1993, Geoff Kuenning, Granada Hills, CA All rights reserved.

Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:

1. Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.

2. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.

3. All modifications to the source code must be clearly marked as such. Binary redistributions based on modified source code must be clearly marked as modified versions in the documentation and/or other materials provided with the distribution.

4. All advertising materials mentioning features or use of this software must display the following acknowledgment:

This product includes software developed by Geoff Kuenning and other unpaid contributors.

5. The name of Geoff Kuenning may not be used to endorse or promote products derived from this software without specific prior written permission.

THIS SOFTWARE IS PROVIDED BY GEOFF KUENNING AND CONTRIBUTORS “AS IS” AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICU-LAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL GEOFF KUENNING OR CON-TRIBUTORS BE LIABLE FOR ANY DIRECT, INDIRECT, INCIDENTAL, SPECIAL, EXEM-PLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO, PRO-CUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR PROFITS; OR BUSINESS INTERRUPTION) HOWEVER CAUSED AND ON ANY THEORY OF LIA-BILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING NEGLI-GENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.

F.3 The Blitz++ Artistic License

(with thanks and apologies to authors of the Perl Artistic License)

Preamble

The intent of this document is to state the conditions under which Blitz++ may be copied, such that the authors maintains some semblance of artistic control over the development of the package, while giving the users of the package the right to use and distribute Blitz++ in a more-or-less customary fashion, plus the right to make reasonable modifications.

Definitions

Library refers to the collection of files distributed by the Copyright Holder, and derivatives of that collection of files created through textual modification.

Standard Version refers to such a Library if it has not been modified, or has been modified in accordance with the wishes of the Copyright Holder as specified below.

Copyright Holder’ is whoever is named in the copyright or copyrights for the package.

You is you, if you’re thinking about copying, modifying or distributing this Library.

Freely Available means that no fee is charged for the item. It also means that recipients of the item may redistribute it under the same conditions they received it.

‘Reasonable copying fee’ is whatever you can justify on the basis of media cost, duplication charges, time of people involved, and so on. (You will not be required to justify it to the Copyright Holder, but only to the computing community at large as a market that must bear the fee.)

1. You may make and give away verbatim copies of the Standard Version of this Library without restriction, provided that you duplicate all of the original copyright notices, this license, and associated disclaimers.

2. The Standard Version of the Library may be distributed as part of a collection of software, provided no more than a reasonable copying fee is charged for the software collection.

3. You may apply bug fixes, portability fixes and other modifications derived from the Public Domain or from the Copyright Holder. A Library modified in such a way shall still be considered the Standard Version.

4. You may otherwise modify your copy of this Library in any way, provided that you insert a prominent notice in each changed file stating how and when you changed that file, and provided that you do at least ONE of the following:

a. place your modifications in the Public Domain or otherwise make them Freely Available, such as by posting said modifications to the Blitz++ development list, and allowing the Copyright Holder to include your modifications in the Standard Version of the Library.

b. use the modified Library only within your corporation or organization.

c. make other distribution arrangements with the Copyright Holder.

5. You may distribute programs which use this Library in object code or executable form without restriction.

6. Any object code generated as a result of using this Library does not fall under the copyright of this Library, but belongs to whomever generated it, and may be sold commercially.

7. The name of the Copyright Holder or the Library may not be used to endorse or promote products derived from this software without specific prior written permission.

8. THIS PACKAGE IS PROVIDED AS IS AND WITHOUT ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, WITHOUT LIMITATION, THE IMPLIED WARRANTIES OF MERCHANTIBILITY AND FITNESS FOR A PARTICULAR PURPOSE.